OpenStudy (anonymous):

Let a = (-8, 0, 4) and b = (-5, 0, 0) be vectors. Find the scalar, vector, and orthogonal projections of b onto a. Scalar Projection: ??? Vector Projection: ( ? , ? , ? ) Orthogonal Projection: ( ? , ? , ? )

OpenStudy (anonymous):

scalar proj: a dot b / magnitude of a = 40/√80 = 4.47 vector proj : (a dot b / magnitude of a squared) * a = 40/80 * (-8, 0, 4) = (-4, 0, 2) orth proj: weird name for the cross product of a and b but i assume thats what you are asked for, a X b = (0, 20, 0)

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

sure thing